# Deconstructing Obfuscation: A four-dimensional framework for evaluating Large Language Models assembly code deobfuscation capabilities

### v1

@@ 4 Discussion
Our systematic evaluation of several state-of-the-art LLMs across multiple obfuscation techniques reveals significant performance variations, from autonomous deobfuscation to complete failure, depending on the technique employed. Rather than attributing these differences to singular factors like model size or architecture, we propose a theoretical framework based on four distinct yet interconnected dimensions: Reasoning Depth (analytical processing of program logic), Pattern Recognition (structural identification amid obfuscation), Noise Filtering (distinguishing essential logic from deceptive constructs), and Context Integration (maintaining coherence across fragmented code). This framework provides definitive explanatory power for our results and establishes predictive value for future advances. The following subsections explore each dimension through established program analysis theory, analyze cybersecurity implications, address study limitations, and identify promising research directions for enhancing AI-assisted reverse engineering.
@@ 4.1 Theoretical Framework for LLM Deobfuscation
Context Integration refers to a model’s ability to maintain coherence across logically related but physically disconnected code segments. This dimension parallels control flow graph analysis [54] and leverages transformer architecture’s self-attention mechanisms [4], which capture long-range dependencies across token sequences. Our experiments revealed varied capabilities: GPT-4.5, GPT-o1, and Grok 3 demonstrated superior integration, successfully reconstructing relationships between computations at address 0x1585 with condition input & 3 == 0 (Section 3.3), despite fragmentation through state machines initialized with constant 0x64fd8a96. Claude 3.7 Sonnet correctly determined that jumps to address 0x40169e were unreachable due to opaque predicates but struggled with control flow flattening (Section 3.1). DeepSeek R1 and GPT-3o Mini produced state machine representations with transitions like state = 0x4c3945a8 to state = 0xd87b0953 (Section 3.4.4 and Section 3.1.4) without simplifying to underlying algorithms. When facing combined obfuscation, all models struggled, with even GPT-4.5 misinterpreting intermediate state values at 0x1579 as definitive outputs (Section 3.4.1). We identified a clear relationship between context window size and performance, models with windows exceeding 200,000 tokens demonstrated superior performance on complex tasks.
@@
Pattern Recognition involves identifying structural and computational patterns within obfuscated code, building upon program similarity metrics and clone detection theory [50, 51]. This dimension connects to the "naturalness of software" hypothesis [52], which proposed that programming languages contain statistical regularities enabling probabilistic modeling. Our experiments showed that models like GPT-4.5, GPT-o1, and Grok 3 successfully reconstructed the underlying switch structures despite extensive obfuscation (Section 3.3), identifying complex arithmetic transformations such as (input | 0xbaaad0bf) * (input ^ 2) within heavily obfuscated contexts (Section 3.2). This capability emerged organically from general language modeling rather than specialized program analysis tools. Performance varied significantly between obfuscation techniques - models that excelled at reconstructing control structures often struggled with arithmetic transformations, supporting the findings that statistical models of code can capture both syntactic and semantic patterns [53], with varying degrees of success.
@@
Noise Filtering encompasses a model’s ability to distinguish essential computational logic from obfuscation artifacts, sharing functional similarities with program slicing [48] despite fundamentally different mechanisms. From an information-theoretic perspective [47], LLMs must extract meaningful signal from the increased entropy that obfuscation introduces. Our experiments demonstrated this capability in specific contexts: when analyzing bogus control flow, models like GPT-4.5, GPT-o1, and Claude 3.7 Sonnet correctly identified repetitive predicate checks as non-functional (Section 3.1.2, Section 3.1.3, Section 3.1), simplifying the code to its essential form. However, we observed significant limitations: DeepSeek R1 generated convoluted expressions like ((0xBAAAD0BF | n) ^ (0xE6C98769 & ~n) ^ 2) instead of the simpler (n | 0xBAAAD0BF) * (n ^ 2) (Section 3.2.4). This parallels challenges in adversarial machine learning [49], where obfuscation disrupts the statistical patterns on which LLMs rely, creating noise that confounds even models with strong capabilities in other dimensions.
@@ 4.2 Mapping Framework Dimensions to Empirical Results
Our analysis revealed important dimensional performance patterns that open new valuable research directions. For instance, Claude 3.7 Sonnet demonstrated exceptional Reasoning Depth with bogus control flow (Level 0) (Section 3.1) but showed comparatively weaker Context Integration with control flow flattening (Level 0-1) (Section 3.3.4), despite both techniques requiring sophisticated reasoning capabilities. Similarly, models with large context windows occasionally performed worse than smaller-context models on tasks that seemingly should benefit from extended context. These patterns demonstrate significant interactions between dimensional capabilities and specific implementation details of obfuscation techniques that our current framework does not fully capture. Future research should specifically investigate these boundary cases to refine our understanding of how dimensional capabilities interact and potentially identify additional factors that influence deobfuscation performance.
@@ 4.2.1 Taxonomy of LLM Deobfuscation Errors
These errors involve failing to recognize that certain conditions are always true or false (invariant), relating directly to limitations in the Reasoning Depth dimension of our framework. When analyzing bogus control flow, GPT-4o misinterpreted the opaque predicate (var1 * (var1 - 1)) & 1 == 0 as potentially false (Section 3.1.1), despite the mathematical property that the product of consecutive integers is always even. This misinterpretation led to incorrect control flow analysis, with the model treating the unreachable jump to address 0x40169e as a viable execution path. Similarly to the challenges faced in formal verification [46], these errors demonstrate the difficulty in symbolically reasoning about invariant properties. This error was prevalent in the analysis of bogus control flow obfuscation, appearing in 4 of 7 models (GPT-4o, GPT-3o Mini, Grok 2, DeepSeek R1 in Section 3.1). In contrast, models with enhanced reasoning capabilities, such as Claude 3.7 Sonnet, GPT-4.5, and Grok 3 (in Section 3.1), correctly identified the predicate’s invariant property. Beyond its frequency, this error has practical consequences, such as misidentification of potentially exploitable code paths in vulnerability analysis.
@@
These errors occur when a model correctly identifies computational components but incorrectly maps them to control structures, reflecting limitations in both the Pattern Recognition and Context Integration dimensions of our framework. When analyzing CFF - Grok 2 demonstrated this error type, correctly identifying four arithmetic transformations but assigning three to incorrect conditional paths. Specifically, it matched (input + 0xbaaad0bf) * (input & 5) to input & 3 == 0 instead of input & 3 == 3 (Section 3.3.1). This creates deobfuscated code that preserves computational elements but misimplements their control dependencies, meaning that the conditions governing their execution are wrong, resulting in runtime behavior fundamentally different from the original despite structural similarity. This type of error presents a particularly challenging issue for code analysis applications, where incorrect mappings may pass superficial validation, yet introduce subtle semantic errors that fundamentally alter program behavior[53]. We observed this error in 3 of 7 models (Grok 2, DeepSeek R1, GPT-4o in Section 3.3), while models with stronger context integration capabilities, such as larger context windows or advanced attention mechanisms (e.g., GPT-4.5 and Claude 3.7 Sonnet in Section 3.3), correctly reconstructed these relationships, particularly when dealing with less complex obfuscation techniques.
@@
This error type, primarily reflecting limitations in our framework’s Noise Filtering and Pattern Recognition dimensions, occurs when models incorrectly reconstruct the fundamental control flow structure of the original code. A common manifestation we observed was the incorrect inference of iterative structures (loops, state machines) where only single-pass logic exists. When analyzing bogus control flow obfuscation, DeepSeek R1 and GPT-3o Mini misinterpreted repetitive predicate checks (deliberately inserted as obfuscation artifacts) as indicators of functional looping behavior. For example, DeepSeek R1 generated for_n range(10) constructs in its deobfuscated output, despite the original assembly implementing a non-iterative execution path (Section 3.1.6). Similarly, when facing combined obfuscation techniques, ChatGPT-pro-o1 and Grok 3 reconstructed the code as a looping state machine (e.g., while (1) switch (state) ... (Section 3.4.2 and Section 3.4.3)) rather than identifying the single-pass switch statement at address 0x156b. While this error resembles hallucination phenomena in other LLM applications [55], our evidence demonstrates it primarily arises from misclassifying obfuscation patterns. These patterns, such as control flow flattening, transform direct branches into state-machine-like constructs to mislead static analysis tools. This misinterpretation carries significant security implications, as it can fundamentally alter an analyst’s understanding of program execution, potentially obscuring vulnerabilities or introducing nonexistent execution paths. These consequences directly impact vulnerability assessment procedures, where accurate control flow reconstruction is essential to identify potentially exploitable paths [56, 57], particularly in contexts where automated tools supplement human analysis. We observed this error in 5 out of 7 models (DeepSeek R1, GPT-3o Mini, ChatGPT-pro-o1, Grok 3, Claude 3.7 Sonnet), highlighting a widespread difficulty in separating obfuscation from actual control flow.
@@
These errors involve failing to correctly reconstruct precise arithmetic and bitwise operations from obfuscated assembly, directly reflecting limitations in the Pattern Recognition dimension of our framework. All seven models exhibited this error type when confronting instruction substitution techniques, though with varying severity. Our analysis documented systematic pattern failures: ChatGPT-4o (Section 3.2.1) replaced the operation (input | 0xBAAAD0BF) * (input ^ 2) with entirely incorrect expressions containing unrelated constants and inappropriate operations. GPT-4.5 demonstrated a different failure mode (Section 3.2.2), preserving core operations but introducing unnecessary complexity by adding extraneous bitwise operations and constants to what should be simple expressions. For example, instead of the straightforward operation present in the original code, it generated (( input ^ 0xBAAAD0BF ^ 0xE6C98769) + 2) * (input | 2) . DeepSeek R1 (Section 3.2.4) produced particularly convoluted transformations with multiple nested operations and unrelated constants. Models adopted different strategies when handling obfuscated arithmetic operations. Claude 3.7 Sonnet (Section 3.2.5) attempted mathematical abstraction, deriving patterns like (-3) * (x ^ 2) for even-indexed cases while ignoring specific constants. GPT-4.5 preserved core operations but introduced unnecessary complexity. Grok 3(Section 3.2.3) achieved partial accuracy, correctly reconstructing some operations while misrepresenting others. DeepSeek R1 produced the most convoluted transformations with multiple nested operations and fabricated constants. These diverse approaches, each with distinct trade-offs between abstraction and fidelity, demonstrate the fundamental challenge LLMs face when attempting to simplify deliberately obfuscated arithmetic. The consistency of these errors across all models when faced with instruction substitution, a technique specifically designed to obscure individual operations, suggests a fundamental limitation in the ability of current LLMs to recognize equivalent computational expressions, particularly when obfuscation deliberately transforms simple operations into more complex forms through mathematical identities (all examples from Section 3.2).
@@
These errors involve incorrectly handling, identifying, or fabricating literal values in deobfuscated code, revealing limitations in both Pattern Recognition and Context Integration dimensions. We identified three distinct manifestations with varying severity. First, subtle transcription errors occurred when models processed large hexadecimal constants, exemplified by Claude 3.7 Sonnet’s (Section 3.2.5) misreading of 0xBAAAD0BF as 0xBAAAAD0BF (inserting an extra ’A’) during instruction substitution analysis. Second, models introduced mathematically unrelated constants, as seen when ChatGPT-4.5 (Section 3.2.2) generated expressions containing values like 0xE6C98769 and 0x8ABD1CD5 — constants not present in the original assembly but erroneously incorporated into arithmetic operations. Most concerning was the third category: complete fabrication of well-known "hexspeak" values. When analyzing combined obfuscation techniques, DeepSeek R1 generated deobfuscated code containing entirely fictional constants (Section 3.4.4) like 0xBAADF00D, 0xDEADBEEF, and 0xCAFEBABE—recognizable placeholder values commonly used by developers but completely absent from the original assembly. This fabrication suggests a problematic pattern-matching behavior where models revert to familiar training examples rather than faithfully representing the analyzed code. From a security perspective, these errors are particularly troubling when analyzing cryptographic implementations or any code where precise constant values directly impact functionality and correctness, such as protocol implementations, hash functions, or file format parsers. We identified constant propagation errors in 3 of 7 models (Claude 3.7 Sonnet, DeepSeek R1, GPT-4.5 in (Section 3.2)), indicating that even advanced LLMs struggle with accurate representation of symbolic information.
@@ 4.2.2 Obfuscation Technique Resistance Model
Bogus control flow obfuscation primarily challenges the Reasoning Depth dimension through opaque predicates and unreachable code paths. Our experiments revealed clear performance stratification: top performers (Claude 3.7 Sonnet, GPT-4.5, Grok 3) required minimal or no guidance (Levels 0-2), while others needed expert intervention or failed entirely. This stratification directly correlates with each model’s ability to perform mathematical reasoning about invariant conditions. Models with stronger reasoning capabilities performed consistently better (Section 3), demonstrating that this obfuscation technique, traditionally effective against static analysis tools, offers limited protection against advanced LLMs with enhanced reasoning capabilities.
@@
Control flow flattening primarily challenges the Context Integration dimension by fragmenting logically connected code segments across a state machine structure. This technique demonstrated intermediate resistance, with only two models achieving Level 0-1 success. Most models either required substantial guidance or failed entirely, frequently producing complex state machine representations that failed to simplify to the underlying algorithm. Effective deobfuscation requires both strong pattern recognition to identify computational elements and robust context integration to correctly reconstruct their relationships—a combination present in fewer models than those capable of handling bogus control flow.

### v2

@@ 4 Discussion
Our systematic evaluation of several state-of-the-art LLMs across multiple obfuscation techniques reveals significant performance variations, from autonomous deobfuscation to complete failure, depending on the technique employed. Rather than attributing these differences to singular factors like model size or architecture, we propose a theoretical framework based on four distinct yet interconnected dimensions: Reasoning Depth (analytical processing of program logic), Pattern Recognition (structural identification amid obfuscation), Noise Filtering (distinguishing essential logic from deceptive constructs), and Context Integration (maintaining coherence across fragmented code). This framework provides substantial explanatory power for our results and establishes predictive value for future advances. The following subsections explore each dimension through established program analysis theory, analyze cybersecurity implications, address study limitations, and identify promising research directions for enhancing AI-assisted reverse engineering.
@@ 4.1 Theoretical Framework for LLM Deobfuscation
Context Integration refers to a model’s ability to maintain coherence across logically related but physically disconnected code segments. This dimension parallels control flow graph analysis [54] and leverages transformer architecture’s self-attention mechanisms [4], which capture long-range dependencies across token sequences. Our experiments revealed varied capabilities: GPT-4.5, GPT-Pro-o1, and Grok 3 demonstrated superior integration, successfully reconstructing relationships between computations at address 0x1585 with condition input & 3 == 0 (Section 3.3), despite fragmentation through state machines initialized with constant 0x64fd8a96. DeepSeek R1 and GPT-3o Mini produced state machine representations with transitions like state = 0x4c3945a8 to state = 0xd87b0953 (Section 3.4.4 and Section 3.1.4) without simplifying to underlying algorithms. When facing combined obfuscation, all models struggled, with even GPT-4.5 misinterpreting intermediate state values at 0x1579 as definitive outputs (Section 3.4.1). We identified a clear relationship between context window size and performance, models with windows exceeding 200,000 tokens demonstrated superior performance on complex tasks.
@@
Pattern Recognition involves identifying structural and computational patterns within obfuscated code, building upon program similarity metrics and clone detection theory [50, 51]. This dimension connects to the "naturalness of software" hypothesis [52], which proposed that programming languages contain statistical regularities enabling probabilistic modeling. Our experiments showed that models like GPT-4.5, GPT-Pro-o1, and Grok 3 successfully reconstructed the underlying switch structures despite extensive obfuscation (Section 3.3), identifying complex arithmetic transformations such as (input | 0xbaaad0bf) * (input ^ 2) within heavily obfuscated contexts (Section 3.2). This capability emerged organically from general language modeling rather than specialized program analysis tools. Performance varied significantly between obfuscation techniques - models that excelled at reconstructing control structures often struggled with arithmetic transformations, supporting the findings that statistical models of code can capture both syntactic and semantic patterns [53], with varying degrees of success.
@@
Noise Filtering encompasses a model’s ability to distinguish essential computational logic from obfuscation artifacts, sharing functional similarities with program slicing [48] despite fundamentally different mechanisms. From an information-theoretic perspective [47], LLMs must extract meaningful signal from the increased entropy that obfuscation introduces. Our experiments demonstrated this capability in specific contexts: when analyzing bogus control flow, models like GPT-4.5, GPT-Pro-o1, and Claude 3.7 Sonnet correctly identified repetitive predicate checks as non-functional (Section 3.1.2, Section 3.1.3, Section 3.1), simplifying the code to its essential form. However, we observed significant limitations: DeepSeek R1 generated convoluted expressions like ((0xBAAAD0BF | n) ^ (0xE6C98769 & ~n) ^ 2) instead of the simpler (n | 0xBAAAD0BF) * (n ^ 2) (Section 3.2.4). This parallels challenges in adversarial machine learning [49], where obfuscation disrupts the statistical patterns on which LLMs rely, creating noise that confounds even models with strong capabilities in other dimensions.
@@ 4.2 Mapping Framework Dimensions to Empirical Results
Our analysis revealed important dimensional performance patterns that open new valuable research directions. For instance, Claude 3.7 Sonnet demonstrated exceptional Reasoning Depth with bogus control flow (Level 0) (Section 3.1) and showed similar Context Integration with control flow flattening (Level 0-1) (Section 3.3.4), despite both techniques requiring sophisticated reasoning capabilities. Similarly, models with large context windows occasionally performed worse than smaller-context models on tasks that seemingly should benefit from extended context. These patterns demonstrate significant interactions between dimensional capabilities and specific implementation details of obfuscation techniques that our current framework does not fully capture. Future research should specifically investigate these boundary cases to refine our understanding of how dimensional capabilities interact and potentially identify additional factors that influence deobfuscation performance.
@@ 4.2.1 Taxonomy of LLM Deobfuscation Errors
These errors involve failing to recognize that certain conditions are always true or false (invariant), relating directly to limitations in the Reasoning Depth dimension of our framework. When analyzing bogus control flow, GPT-4o misinterpreted the opaque predicate (var1 * (var1 - 1)) & 1 == 0 as potentially false (Section 3.1.1), despite the mathematical property that the product of consecutive integers is always even. This misinterpretation led to incorrect control flow analysis, with the model treating the unreachable jump to address 0x40169e as a viable execution path. Similarly to the challenges faced in formal verification [46], these errors demonstrate the difficulty in symbolically reasoning about invariant properties. This error was prevalent in the analysis of bogus control flow obfuscation, appearing in 3 of 8 models (GPT-4o, GPT-3o Mini, DeepSeek R1 in Section 3.1). In contrast, models with enhanced reasoning capabilities, such as Claude 3.7 Sonnet, GPT-4.5, and Grok 3 (in Section 3.1), correctly identified the predicate’s invariant property. Beyond its frequency, this error has practical consequences, such as misidentification of potentially exploitable code paths in vulnerability analysis.
@@
These errors occur when a model correctly identifies computational components but incorrectly maps them to control structures, reflecting limitations in both the Pattern Recognition and Context Integration dimensions of our framework. When analyzing CFF - Grok 2 demonstrated this error type, correctly identifying four arithmetic transformations but assigning three to incorrect conditional paths. Specifically, it matched (input + 0xbaaad0bf) * (input & 5) to input & 3 == 0 instead of input & 3 == 3 (Section 3.3.1). This creates deobfuscated code that preserves computational elements but misimplements their control dependencies, meaning that the conditions governing their execution are wrong, resulting in runtime behavior fundamentally different from the original despite structural similarity. This type of error presents a particularly challenging issue for code analysis applications, where incorrect mappings may pass superficial validation, yet introduce subtle semantic errors that fundamentally alter program behavior[53]. This type of error was observed during the deobfuscation process, with varying outcomes depending on model capabilities. While some models like Grok 2 initially exhibited structural mapping errors but recovered with guidance (Section 3.3.1), models with stronger context integration capabilities (e.g., GPT-4.5 and Claude 3.7 Sonnet) correctly reconstructed these relationships from the outset, achieving autonomous deobfuscation (Level 0-1). This demonstrates that structural mapping represents a recoverable error for capable models rather than a fundamental limitation.
@@
This error type, primarily reflecting limitations in our framework’s Noise Filtering and Pattern Recognition dimensions, occurs when models incorrectly reconstruct the fundamental control flow structure of the original code. A common manifestation we observed was the incorrect inference of iterative structures (loops, state machines) where only single-pass logic exists. When analyzing bogus control flow obfuscation, DeepSeek R1 and GPT-3o Mini misinterpreted repetitive predicate checks (deliberately inserted as obfuscation artifacts) as indicators of functional looping behavior. For example, DeepSeek R1 generated for_n range(10) constructs in its deobfuscated output, despite the original assembly implementing a non-iterative execution path (Section 3.1.6). Similarly, when facing combined obfuscation techniques, ChatGPT-pro-o1 and Grok 3 reconstructed the code as a looping state machine (e.g., while (1) switch (state) ... (Section 3.4.2 and Section 3.4.3)) rather than identifying the single-pass switch statement at address 0x156b. While this error resembles hallucination phenomena in other LLM applications [55], our evidence demonstrates it primarily arises from misclassifying obfuscation patterns. These patterns, such as control flow flattening, transform direct branches into state-machine-like constructs to mislead static analysis tools. This misinterpretation carries significant security implications, as it can fundamentally alter an analyst’s understanding of program execution, potentially obscuring vulnerabilities or introducing nonexistent execution paths. These consequences directly impact vulnerability assessment procedures, where accurate control flow reconstruction is essential to identify potentially exploitable paths [56, 57], particularly in contexts where automated tools supplement human analysis. We observed this error in 3 out of 8 models (DeepSeek R1, GPT-3o Mini, ChatGPT-pro-o1), highlighting a widespread difficulty in separating obfuscation from actual control flow.
@@
These errors involve failing to correctly reconstruct precise arithmetic and bitwise operations from obfuscated assembly, directly reflecting limitations in the Pattern Recognition dimension of our framework. All five models that completed instruction substitution evaluation exhibited this error type, though with varying severity. Our analysis documented systematic pattern failures: ChatGPT-4o (Section 3.2.1) replaced the operation (input | 0xBAAAD0BF) * (input ^ 2) with entirely incorrect expressions containing unrelated constants and inappropriate operations. GPT-4.5 demonstrated a different failure mode (Section 3.2.2), preserving core operations but introducing unnecessary complexity by adding extraneous bitwise operations and constants to what should be simple expressions. For example, instead of the straightforward operation present in the original code, it generated (( input ^ 0xBAAAD0BF ^ 0xE6C98769) + 2) * (input | 2) . DeepSeek R1 (Section 3.2.4) produced particularly convoluted transformations with multiple nested operations and unrelated constants. Models adopted different strategies when handling obfuscated arithmetic operations. Claude 3.7 Sonnet (Section 3.2.5) attempted mathematical abstraction, deriving patterns like (-3) * (x ^ 2) for even-indexed cases while ignoring specific constants. GPT-4.5 preserved core operations but introduced unnecessary complexity. Grok 3(Section 3.2.3) achieved partial accuracy, correctly reconstructing some operations while misrepresenting others. DeepSeek R1 produced the most convoluted transformations with multiple nested operations and fabricated constants. These diverse approaches, each with distinct trade-offs between abstraction and fidelity, demonstrate the fundamental challenge LLMs face when attempting to simplify deliberately obfuscated arithmetic. The consistency of these errors across all models when faced with instruction substitution, a technique specifically designed to obscure individual operations, suggests a fundamental limitation in the ability of current LLMs to recognize equivalent computational expressions, particularly when obfuscation deliberately transforms simple operations into more complex forms through mathematical identities (all examples from Section 3.2).
@@
These errors involve incorrectly handling, identifying, or fabricating literal values in deobfuscated code, revealing limitations in both Pattern Recognition and Context Integration dimensions. We identified three distinct manifestations with varying severity. First, subtle transcription errors occurred when models processed large hexadecimal constants, exemplified by Claude 3.7 Sonnet’s (Section 3.2.5) misreading of 0xBAAAD0BF as 0xBAAAAD0BF (inserting an extra ’A’) during instruction substitution analysis. Second, models introduced mathematically unrelated constants, as seen when ChatGPT-4.5 (Section 3.2.2) generated expressions containing values like 0xE6C98769 and 0x8ABD1CD5 — constants not present in the original assembly but erroneously incorporated into arithmetic operations. Most concerning was the third category: complete fabrication of well-known "hexspeak" values. When analyzing combined obfuscation techniques, DeepSeek R1 generated deobfuscated code containing entirely fictional constants (Section 3.4.4) like 0xBAADF00D, 0xDEADBEEF, and 0xCAFEBABE—recognizable placeholder values commonly used by developers but completely absent from the original assembly. This fabrication suggests a problematic pattern-matching behavior where models revert to familiar training examples rather than faithfully representing the analyzed code. From a security perspective, these errors are particularly troubling when analyzing cryptographic implementations or any code where precise constant values directly impact functionality and correctness, such as protocol implementations, hash functions, or file format parsers. We identified constant propagation errors in 3 of 8 models (Claude 3.7 Sonnet, DeepSeek R1, GPT-4.5 in (Section 3.2)), indicating that even advanced LLMs struggle with accurate representation of symbolic information.
@@ 4.2.2 Obfuscation Technique Resistance Model
Control flow flattening primarily challenges the Context Integration dimension by fragmenting logically connected code segments across a state machine structure. Our experiments revealed excellent performance across most models: several top performers (GPT-4.5, GPT-Pro-o1, Grok 3) achieved autonomous deobfuscation (Level 0), while most others required minimal guidance (Levels 0-2). Models with larger context windows and advanced attention mechanisms performed particularly well (Section 3.3), suggesting that this obfuscation technique, while effective against traditional static analysis tools, offers limited protection against LLMs with strong context integration capabilities.
@@
Bogus control flow obfuscation primarily challenges the Reasoning Depth dimension through opaque predicates and unreachable code paths. This technique demonstrated intermediate resistance, with clear performance stratification across models: only one model achieved autonomous deobfuscation (Claude 3.7 Sonnet at Level 0), while two others required minimal guidance (Grok 3 at Level 1, GPT-4.5 at Levels 1-2). However, the remaining models needed substantial expert intervention or failed entirely (GPT-Pro-o1 at Level 3, GPT-4o at Level 4, GPT-3o Mini and DeepSeekR1 at Levels 4-5). This stratification directly correlates with each model’s ability to perform mathematical reasoning about invariant conditions (Section 3.1). The mixed performance demonstrates that while advanced models with enhanced reasoning capabilities can overcome this obfuscation, it still presents significant challenges for models with weaker reasoning depth, making it moderately effective against current LLM-based analysis.